# From Wavefunctional Entanglement to Entangled Wavefunctional Degrees of Freedom

### v1

@@ Abstract
The question of whether entanglement between photons is equivalent to entanglement between their characteristic field modes—specifically, the single-particle wavefunctions that are composed and superposed to describe identical particles—is a key, open problem concerning multi-partite optical degrees of freedom, and has profound implications for topics ranging from quantum foundations to quantum computation. In this Letter, I offer a fresh, deeper, physical insight into this subtle, albeit enduring, issue by describing a situation in which entangling interactions between optical modes—namely, the wavefunctions—can be distilled into genuine entanglement between the physical properties of the photons—which are the wavefunctional degrees of freedom. This theoretical observation also highlights the salience of the measurement context—especially, of clearly disambiguating between the choice of the quantum subsystem and the decision to measure an observable along a particular axis of measurement—while quantifying and transforming quantum optical entanglement. This understanding might be applied to formulate a new class of protocols for distilling quantum magic from contextually and nonlocally entangled photons within inseparable field modes.
Introduction.—Quantum entanglement and quantum nonlocality [1, 2, 3, 4, 5, 6, 7, 8] are fundamental features of the quantum mechanical description of many-particle systems, and play defining roles in foundations of quantum physics [9, 10, 11, 12]; quantum information science and engineering [13, 14, 15, 16]; atomic [17, 18, 19, 20], molecular [21], optical [22, 23, 24, 25, 26, 27], condensed matter [28, 29], and high-energy physics [30]; as well as in cosmology [31, 32, 33, 34]. While the earliest intimations of the notion of quantum nonlocality are traceable to the 1927 Solvay Conference [*[See, e.g., ][forahistoryofthoughtonentanglement, nonlocality, andincompleteness.]Bacciagaluppi_Crull_2024, 36, 37], in 1935, Einstein, Podolsky, and Rosen (E.P.R.) [1] formally introduced the concept of quantum entanglement in the course of devising a Gedankenexperiment to clearly show a conflict between: the quantum mechanical postulate that the wavefunction provides a complete description of the physical reality of a quantum system; and the absence of simultaneous elements of physical reality corresponding to values of mutually complementary physical quantities that are described by non-commuting operators. Specifically, E.P.R. argued that either or is true, but not both. While, in the standard quantum theory, is true and is false, E.P.R. were able to demonstrate that the situation in which the coordinates, as well as the momenta of two particles are entangled, paradoxically suggests that is false, but is true.
@@
Background.—Guided by Bohm’s reformulation of the E.P.R. paradox [39] in terms of discrete-valued, dichotomous observables, namely, spin angular momenta, Bell correctly realized that the above-described conflict is actually one between the irrevocable indeterminism of quantum mechanics and E.P.R.’s presupposed, cherished notion of local, classical realism [40, *bell2004speakablepaper2]. Simply put, situations, such as the E.P.R.B. (E.P.R.-Bohm) entanglement in which there are perfect correlations—or anti-correlations—between the outcomes of measurements performed upon two causally disconnected, space-like separated particles suggest that there are two opposing descriptions of quantum reality, namely, either: physical reality is non-local, measurement outcomes are inescapably probabilistic, and the wavefunction description of physical reality is complete; or physical reality is local, measurement outcomes are predetermined by hidden elements of reality, and therefore the wavefunction description of physical reality is incomplete, which is the crux of the E.P.R. argument [42, *werner2014comment, *wiseman2014two].
@@
Notably, Bell resolved the E.P.R. paradox by formulating empirically verifiable inequalities [40, *bell2004speakablepaper2] that quantified the difference between and above. Bell’s inequalities bound from above the strengths of correlations between the outcomes of measurements on two quantum systems in situations in which each of such systems has its own, individual properties, as is suggested by local realist theories that are formulated in the spirit of the E.P.R. viewpoint, like above.
@@
A series of exceedingly elegant and increasingly loophole-free experiments, which were carried out between 1972 and the present time—see, e.g., Refs. [22, 23, 24, 25, 26, 45, 46, 47, 48, 49, 50], and references therein—conclusively demonstrated an explicit violation of refined versions of Bell’s inequalities [51, 52, 53], such as the Bell-C.H.S.H. inequalities [51]. These experiments are enormously significant, as they decisively rule out any conceivable, deterministic, local realist, hidden-variables theory [54, *bell2004speakablepaper1] that attribute the statistical nature of measurement outcomes to the underlying statistical distributions of hidden, supplementary elements of physical reality, analogous to classical probability distributions of classical random variables.
@@
Strikingly, Greenberger, Horne, and Zeilinger (G.H.Z.) extended Bell’s results by showing that—under special circumstances—the definite, certain predictions about measurement outcomes by standard quantum theory directly conflict the corresponding predictions by local realist theories, as opposed to a mere statistical contradiction involving inequalities based on correlation functions [56, 57]. Such situations are realized by probing specific kinds of multi-particle entangled states, known as the G.H.Z. states [58, 59, 27]. -particle G.H.Z. states, where , are commonly used in optical quantum computing schemes and protocols [60, 61, 62, 63, 64, 65, 66, 67, 68, 69] that, fundamentally, derive their quantum advantage from the intrinsic quantum nonlocality of such highly entangled states [70, 60]. More broadly, scalable, quantum computational advantage has been shown to originate from quantum contextuality, which is a generalized notion of quantum nonlocality [71]. Moreover, the quantum mechanical predictions for these multi-particle entangled states are completely independent of the relative arrangements of the measuring apparatuses in space and time.
@@
Motivation for this work.—The quantum mechanical systems of choice in a majority of the above-described experiments were optical, which raised an accompanying, profound question: (Qu1) Is entanglement between modes of light equivalent to entanglement between individual particles of light? (see, e.g., Refs. [72, 73, 74] for a detailed discussion on the history of, and related work on this key problem.) Resolution of this quantum foundational issue also has significant implications for quantum information processing [75] and quantum-enhanced sensing [76] with highly entangled states of light.
@@
Consequently, one may pose the question, which seems both fundamentally, as well as practically important: (Qu3) Is it possible to transfer the inter-modal entanglement—which is relatively simpler to achieve—to inter-particle entanglement—which is particularly challenging to realize—as deterministically as possible? The purpose of this Letter is to show that, indeed, wavefunctional 111Throughout this Letter, the term wavefunctional is used as the adjectival form of wavefunction, and is unrelated to a wave functional that plays the role of a wavefunction in a quantized theory of wave fields [93]. entanglement—that is, entanglement between mathematically inseparable quantum wavefunctions or optical field modes—can be transformed into direct entangling interactions between optical photons that are described by such modes. An additional step that entails making a measurement on an ancillary photon appears to be irreplaceable. Figure 1 summarizes the problem that this Letter addresses, whereas Fig. 2 furnishes a proposed solution.
Disambiguating between choices of subsystems and of observables.—Before tackling question (Qu3), let us consider experiments of the kind that were pioneered by Aspect et al. [23, 25, 24] for demonstrating violations of Bell’s inequalities, so as to illustrate the role played by the relative context of the quantum measuring arrangements—specifically, that of local basis transformations on individual subsystems—in determining the specific definitions—or identities—of subsystems and structures of observables.
@@
where A and B represent the two quantum subsystems—which are the two distinguishable photons—and H and V represent the horizontal and vertical orientations of the linear polarization, respectively. As is described in the accompanying end matter (also see Fig. 3), the above two-particle entangled state can be unitarily transformed to the one below, by performing individual basis transformations upon each of the two subsystems:
where the single-particle states, and are obtained by unitarily transforming and , respectively, for a given orientation angle. Performing distinct, local, unitary transformations—albeit of the same kind, namely, rotations by different angles—on the two causally disconnected subsystems is valid, since can be used to recover all the well-known expressions for the single- and two-particle, detection probabilities; the coefficients of correlation of linear polarization; as well as the Bell-C.H.S.H. inequalities. The entropies of entanglement of and are identical, and constant, regardless of the relative orientation, .
Remarkably, a similar situation can be orchestrated for entanglement between modes that describe neutral, bosonic atoms; the entangled property is the linear momentum of such atoms [77]. Four output modes are interfered two-by-two in two distinct spatial locations—with individually controllable phases, and —so as to demonstrate Bell nonlocality. The entropies of entanglement are also independent of the relative output phase, , which plays the role of above. A key assumption of this work is optical modes can be treated like any other bosonic mode, e.g., describing cold atoms.
Intriguingly, for modes, which are in photon occupation number entangled relationships with each other, we encounter this issue: The decision to measure the entangled property along, as it were, a particular axis of measurement—or, more precisely, a specifically chosen basis—and the identities—or the definitions—of the individual quantum subsystems, upon which such quantum measurements will be made, are inextricably linked, even though the Bell’s inequalities have been properly reframed for dealing with situations involving continuous, external degrees of freedom [80, 81, 82, 83, 84, 77]. For example, basis transformations concomitantly alter the occupation number—the entangled property—and the definition of the modal creation operator—the identity of the subsystem. For the above-described case of atoms with interfering modes, an additional element of subtlety is introduced by linear momenta—unlike spin angular momenta—along distinct axes commuting with each other. I invite the specialist reader to peruse the end matter for a deeper dive into this issue.
With the above discussion in mind, we are now ready to address (Qu3). The principal findings are summarized in the following principle:
@@
A physicist’s intuitive proof.—While I have given a somewhat more formal proof in the Supplemental Material [85], here, I will outline a simple argument that provides a deeper physical intuition into attacking this formidable problem. Figure 2 schematically summarizes this physical argument. Inspired by Ref. [77], I will model these spatial, bosonic modes as bound states, as opposed to, for example, propagating Gaussian wavepackets. I am specifically assuming that such modes can be mapped onto the bound energy eigenmodes of a simple harmonic oscillator, and of its anharmonically perturbed analogues; a quantum mechanical particle can always, meaningfully, said to be placed in a bound state.
@@
Let us now deploy a quantum mechanical random process that we will utilize to create macroscopic quantum superpositions. To accomplish this, we can introduce an “ancillary” quantum mechanical system, namely, a photon (photon 3 in Fig. 2) propagating through the Young’s double-slit experimental setup [see Fig. 3(a)]. As is well known, the single photon—as it transits through this apparatus—interferes with itself, and, consequently, is described by the superposition: of having gone through the top and bottom slits. Moreover, assume the screen to be photo-conducting, and without a lower half—so as to further emphasize that, unlike in the classical case, the position on the screen, e.g., the upper half, where the photon is detected is uncorrelated with the slit of entry, e.g., the top slit. The screen performs a measurement of the position of the photon. Now—assuming ideal detection and quantum efficiencies—the overall state of the detector and amplifier system, which has access only to photons impinging on the upper half, can be written as the superposition: of having registered and not registered the single photon as a single photo-electron. Assuming that the regions of photo-detection and collection of photo-electrons are sufficiently macroscopic, the probability amplitude, akin to the expression for the transition probability amplitude of a macroscopically delocalized detector atom for this diffraction experiment [87]. describes the spectral representation of the electromagnetic field propagating from the source to the detector.
The probabilistically ejected photo-electron is amplified, thereby producing a photo-current that drives an anharmonic potential controller, which controls the anharmonicity parameter, ; e.g., no photo-current corresponds to the initially harmonic potential. Consequently, the mode wavefunctions are now described as macroscopic superpositions of the energy eigenfunctions of a harmonic oscillator, and their perturbed counterparts corresponding to an anharmonically altered oscillator.

### v2

@@ Abstract
The question of whether entanglement between photons is equivalent to entanglement between their characteristic field modes—specifically, the single-particle wavefunctions that are composed and superposed to describe identical particles—is a key, open problem concerning multi-partite optical degrees of freedom, and has profound implications for topics ranging from quantum foundations to quantum computation. Here, I offer a fresh, deeper, physical insight into this subtle, albeit enduring, issue by describing a situation in which entangling interactions between optical modes—namely, the wavefunctions—can be distilled into genuine entanglement between the physical properties of the photons—which are the wavefunctional degrees of freedom. This theoretical observation also highlights the salience of the measurement context—especially, of clearly disambiguating between the choice of the quantum subsystem and the decision to measure an observable along a particular axis of measurement—while quantifying and transforming quantum optical entanglement. This understanding might be applied to formulate a new class of protocols for distilling quantum magic from contextually and nonlocally entangled photons within inseparable field modes.
I Introduction
Quantum entanglement and quantum nonlocality [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] are fundamental features of the quantum mechanical description of many-particle systems, and play defining roles in foundations of quantum physics [12, 13, 14, 15]; quantum information science and engineering [16, 17, 18, 19]; atomic [20, 21, 22, 23, 24, 25], molecular [26], optical [27, 28, 29, 30, 31, 32], condensed matter [33, 34], and high-energy physics [35]; as well as in cosmology [36, 37, 38, 39]. While the earliest intimations of the notion of quantum nonlocality are traceable to the 1927 Solvay Conference [*[See, forexample, ][forahistoryofthoughtonentanglement, nonlocality, andincompleteness.]Bacciagaluppi_Crull_2024, 41, 42], in 1935, Einstein, Podolsky, and Rosen (E.P.R.) [1] formally introduced the concept of quantum entanglement in the course of devising a Gedankenexperiment to clearly show a conflict between: the quantum mechanical postulate that the wavefunction provides a complete description of the physical reality of a quantum system; and the absence of simultaneous elements of physical reality corresponding to values of mutually complementary physical quantities that are described by non-commuting operators. Specifically, E.P.R. argued that either or is true, but not both. While, in the standard quantum theory, is true and is false, E.P.R. were able to demonstrate that the situation in which the coordinates, as well as the momenta of two particles are entangled, paradoxically suggests that is false, but is true.
@@ I Introduction
Guided by Bohm’s reformulation of the E.P.R. paradox [44] in terms of discrete-valued, dichotomous observables, namely, spin angular momenta, Bell correctly realized that the above-described conflict is actually one between the irrevocable indeterminism of quantum mechanics and E.P.R.’s presupposed, cherished notion of local, classical realism [45, *bell2004speakablepaper2]. Simply put, situations, such as the E.P.R.B. (E.P.R.-Bohm) entanglement in which there are perfect correlations—or anti-correlations—between the outcomes of measurements performed upon two causally disconnected, space-like separated particles suggest that there are two opposing descriptions of quantum reality, namely, either: physical reality is nonlocal, measurement outcomes are inescapably probabilistic, and the wavefunction description of physical reality is complete; or physical reality is local, measurement outcomes are predetermined by hidden elements of reality, and therefore the wavefunction description of physical reality is incomplete, which is the crux of the E.P.R. argument [47, *werner2014comment, *wiseman2014two].
@@
Notably, Bell resolved the E.P.R. paradox by formulating empirically verifiable inequalities [45, *bell2004speakablepaper2] that quantified the difference between and above. Bell’s inequalities bound from above the strengths of correlations between the outcomes of measurements on two quantum systems in situations in which each of such systems has its own, individual properties, as is suggested by local realist theories that are formulated in the spirit of the E.P.R. viewpoint, like above.
@@
A series of exceedingly elegant and increasingly loophole-free experiments, which were carried out between 1972 and the present time—see, for example, Refs. [27, 28, 29, 30, 31, 50, 51, 52, 53, 54, 55], and references therein—conclusively demonstrated an explicit violation of refined versions of Bell’s inequalities [56, 57, 58], such as the Bell-C.H.S.H. inequalities [56]. These experiments are enormously significant, as they decisively rule out any conceivable, deterministic, local realist, hidden-variables theory [59, *bell2004speakablepaper1] that attribute the statistical nature of measurement outcomes to the underlying statistical distributions of hidden, supplementary elements of physical reality, analogous to classical probability distributions of classical random variables.
@@
Strikingly, Greenberger, Horne, and Zeilinger (G.H.Z.) extended Bell’s results by showing that—under special circumstances—the definite, certain predictions about measurement outcomes by standard quantum theory directly conflict the corresponding predictions by local realist theories, as opposed to a mere statistical contradiction involving inequalities based on correlation functions [61, 62]. Such situations are realized by probing specific kinds of multi-particle entangled states, known as the G.H.Z. states [63, 64, 32]. -particle G.H.Z. states, where , are commonly used in optical quantum computing schemes and protocols [65, 66, 67, 68, 69, 70, 71, 72, 73, 74] that, fundamentally, derive their quantum advantage from the intrinsic quantum nonlocality of such highly entangled states [75, 65]. More broadly, multi-particle quantum entanglement and quantum contextuality [59, *bell2004speakablepaper1, 76, 77]—which is a generalized notion of quantum nonlocality and plays an essential role, for example, in magic state distillation protocols [78, 79]—have been collectively suggested [77] as fundamental sources of scalable quantum computational advantage [80, 21, 81] over their classical counterparts. Moreover, the quantum mechanical predictions for these multi-particle entangled states are completely independent of the relative arrangements of the measuring apparatuses in space and time.
II Motivation for this work
@@
The quantum mechanical systems of choice in a majority of the above-described experiments were optical, which raised an accompanying, profound question: (Qu1) Is entanglement between modes of light equivalent to entanglement between individual particles of light? (see, for example, Refs. [82, 83, 84] for a detailed discussion on the history of, and related work on this key problem.) Resolution of this quantum foundational issue also has significant implications for quantum information processing [85] and quantum-enhanced sensing [86] with highly entangled states of light.
@@
Consequently, one might pose the question, which seems both fundamentally, as well as practically important: (Qu3) Is it possible to transfer the inter-modal entanglement—which is relatively simpler to achieve—to inter-particle entanglement—which is particularly challenging to realize—as deterministically as possible? The purpose of this paper is to show that, indeed, wavefunctional 111Throughout this paper, the term wavefunctional is used as the adjectival form of wavefunction, and is unrelated to a wave functional that plays the role of a wavefunction in a quantized theory of wave fields [120]. entanglement—that is, entanglement between mathematically inseparable quantum wavefunctions or optical field modes—can be transformed into direct entangling interactions between optical photons that are described by such modes. An additional step that entails making a measurement on an ancillary photon appears to be irreplaceable. Schematically, Fig. 1 summarizes the problem that this paper addresses, whereas Fig. 2 depicts a proposed solution.
It is important to note that pairs of indistinguishable optical photons—whose angles of linear polarization are entangled—are routinely produced by spontaneous parametric down-conversion (SPDC)-based setups [90]. However, although these sources of entangled photon pairs can be, for example, cascaded [91, 92] to generate three-photon entangled states [92, 93], such highly optically non-linear procedures require complex experimental arrangements [94, 95], and are plagued by low conversion efficiencies, thereby imposing severe requirement on the pump laser. In this paper, I propose an alternative strategy that entails creating mode-mode entangled states—where the modes contain up to photons—and converting such mode-mode entangled states to genuine photon-photon entangled states.
III Disambiguating between choices of subsystems and of observables
Before tackling question (Qu3), let us consider experiments of the kind that were pioneered by Aspect et al. [28, 30, 29] for demonstrating violations of Bell’s inequalities, so as to illustrate the role played by the relative context of the quantum measuring arrangements—specifically, that of local basis transformations on individual subsystems—in determining the specific definitions—or identities—of subsystems and structures of observables.
@@ III Disambiguating between choices of subsystems and of observables
where A and B represent the two quantum subsystems—which are the two distinguishable photons—and H and V represent the horizontal and vertical orientations of the linear polarization, respectively. As is described in Appendix A (see also Fig. 3), the above two-particle entangled state can be unitarily transformed to the one below, by performing individual basis transformations upon each of the two subsystems:
where the single-particle states, and are obtained by unitarily transforming and , respectively, for a given orientation angle. Performing distinct, local, unitary transformations—albeit of the same kind, namely, rotations by different angles—on the two causally disconnected subsystems is valid, since can be used to recover all the well-known expressions for the single- and two-particle, detection probabilities; the coefficients of correlation of linear polarization; as well as the Bell-C.H.S.H. inequalities. The entropies of entanglement of and are identical, and constant, regardless of the relative orientation, . A discussion of the differences between these two representations of the two-particle, entangled state, from a physical point of view, is given in Appendix B.
Remarkably, a similar situation can be orchestrated for entanglement between modes that describe neutral, bosonic atoms; the entangled property is the linear momentum of such atoms [87]. Four output modes are interfered two-by-two in two distinct spatial locations—with individually controllable phases, and —so as to demonstrate Bell nonlocality. The entropies of entanglement are also independent of the relative output phase, , which plays the role of above. A key assumption of this work is optical modes can be treated like any other bosonic mode, for example, describing cold atoms.
Intriguingly, for modes, which are in photon occupation number entangled relationships with each other, we encounter this issue: The decision to measure the entangled property along, as it were, a particular axis of measurement—or, more precisely, a specifically chosen basis—and the identities—or the definitions—of the individual quantum subsystems, upon which such quantum measurements will be made, are inextricably linked, even though the Bell’s inequalities have been properly reframed for dealing with situations involving continuous, external degrees of freedom [96, 97, 98, 99, 100, 87]. For example, basis transformations concomitantly alter the occupation number—the entangled property—and the definition of the modal creation operator—the identity of the subsystem. For the above-described case of atoms with interfering modes, an additional element of subtlety is introduced by linear momenta—unlike spin angular momenta—along distinct axes commuting with each other. I invite the specialist reader to peruse Appendix A for a deeper dive into this issue, and for a discussion of how such ambiguities could be resolved by interferometry.
IV The Equivalence between Wavefunctional and Inter-particle Entanglement
With the above caveats in mind, we are now ready to address (Qu3). The principal findings are summarized in the following principle:
@@ IV The Equivalence between Wavefunctional and Inter-particle Entanglement
While I have given a somewhat more formal proof in Appendix C, here, I will outline a simple argument that provides a deeper physical intuition into attacking this formidable problem. Figure 2 schematically summarizes this physical argument. Inspired by Ref. [87], I will model these spatial, bosonic modes as bound states, as opposed to, for example, propagating Gaussian wavepackets. I am specifically assuming that such modes can be mapped onto the bound energy eigenmodes of a simple harmonic oscillator, and of its anharmonically perturbed analogues; a quantum mechanical particle can always, meaningfully, said to be placed in a bound state.
@@
Let us now deploy a quantum mechanical random process that we will utilize to create macroscopic quantum superpositions. To accomplish this, we can introduce an “ancillary” quantum mechanical system, namely, a photon (photon 3 in Fig. 2) propagating through the Young’s double-slit experimental setup [see Fig. 2(a)]. As is well known, the single photon—as it transits through this apparatus—interferes with itself, and, consequently, is described by the superposition: of having gone through the top and bottom slits. Moreover, assume the screen to be photo-conducting, and without a lower half—so as to further emphasize that, unlike in the classical case, the position on the screen, for example, the upper half, where the photon is detected is uncorrelated with the slit of entry, for example, the top slit. The screen performs a measurement of the position of the photon. Now—assuming ideal detection and quantum efficiencies—the overall state of the detector and amplifier system, which has access only to photons impinging on the upper half, can be written as the superposition: of having registered and not registered the single photon as a single photo-electron. Assuming that the regions of photo-detection and collection of photo-electrons are sufficiently macroscopic, the probability amplitude, akin to the expression for the transition probability amplitude of a macroscopically delocalized detector atom for this diffraction experiment [102]. describes the spectral representation of the electromagnetic field propagating from the source to the detector.
The probabilistically ejected photo-electron is amplified, thereby producing a photo-current that drives an anharmonic potential controller, which controls the anharmonicity parameter, ; for example, no photo-current corresponds to the initially harmonic potential. Consequently, the mode wavefunctions are now described as macroscopic superpositions of the energy eigenfunctions of a harmonic oscillator, and their perturbed counterparts corresponding to an anharmonically altered oscillator.
@@
V A Comparison with Entanglement Swapping
The above-mentioned entangling scheme has several striking similarities to the well-known procedure of entanglement swapping [104, 105, 106, 103], such as projecting the state of two particles onto an entangled state, by making a measurement on an ancillary system. In fact, the fundamental principle of operation underlying both approaches is essentially identical, namely that entanglement is a physical resource that can be transferred between two distinct, non-interacting, uncoupled systems. In what follows, I will briefly review the main idea of entanglement swapping, which I will then compare and contrast with my proposal.
Consider two quantum subsystems, each comprising a pair of initially entangled photons; assume the absence of any direct, physical interactions, or dynamical couplings whatsoever between these subsystems. Now, a suitable kind of projective measurement, such as a joint Bell-state measurement, on two of such photons, each drawn from a distinct initially entangled pair, will project the other two into an entangled state. Moreover, the generation of the results of the Bell-state measurement on two particles heralds that the other two particles have been entangled, thereby allowing, for example, the performance of “event-ready detections” of the entangled particles [103].
Typically, various generalizations of the entanglement swapping scheme entail using pairs of initially entangled photons and projecting the state of two initially unentangled particles onto an entangled state. In contrast, the present scheme transforms the entanglement between optical modes (the wavefunctions) into entanglement between the physical properties of the photons (the wavefunctional degrees of freedom), which are described by these modes. As was pointed out in Ref. [103]: “One could have many different kinds of entanglements to begin with, perform various different measurements, and obtain various kinds of entanglement for the emerging particles.” The present approach, therefore, is a form of entanglement swapping that transforms mode-mode entanglement into particle-particle entanglement.
VI Consequences of Non-Ideal Photodetection Efficiencies
The primary purpose of the simple, ideal Gedankenexperiment, described in Sec. IV, is to point out the possibility of transforming entangled states of modes into entangled states of particles. From a practical point of view, however, the main disadvantage of any realistic scheme, based on this approach, would be its inescapable reliance on successful photodetection of the ancillary photon. Photodetection, typically, has an efficiency, , and, therefore, it is important to consider whether such a scheme would survive realistic photodetection conditions.
An inability to register the ancillary photon that lands on the top half of the screen B (see Fig. 2), due to non-ideal photodetection conditions, can be modeled as an error in forming the state, ; such single-photon-loss errors will cause errors in the probability amplitudes, and , and, consequently, in and . Therefore, the correct output state, will not be formed, and the maximum efficiency of conversion from to will be limited by .
Realistically, the detection screen B can be implemented by a single-photon-sensitive EMCCD (Electron Multiplying Charge-Coupled Device) camera. State-of-the-art EMCCD cameras, typically, have quantum efficiencies (equivalent to the photodetection efficiency, above), that can range up to in the visible wavelengths. Moreover, such cameras have advanced features, such as vacuum thermoelectric cooling and electronic optimization of clock-induced charge, for minimizing spurious photodetection events and the dark noise.
One intriguing idea might to be borrow ideas and concepts from standard quantum error-correction theory to detect and mitigate these ancillary-photon-loss errors. For example, a key finding from recent experiments with neutral atom-based quantum processors is that atom-loss-type leakage errors, if detectable and convertible to erasure errors, are easier to correct than other generic classes of unknown errors [107, 108, 109, 110]. Inspired by these ideas pointing to the feasibility of correcting single-quantum-losses, I wish to propose a quantum error-correcting-based strategy, for mitigating the above-described ancillary-photon-loss errors, that is based on classical control, and that is specific to the experiment in Sec. IV.
Let us imagine that the single-photon source, S is deterministically triggered by a clock signal. The frequency of the clock signal, and, therefore, the single-photon emission rate is kept sufficiently low, such that the time interval between the emissions of two successive photons is greater than the sum of the time it takes the single photon to travel from S to the screen, the time for this photon to be detected and registered, and the time required for a classical, logical AND operation that is described in the next paragraph. Unlike the experiment described in Sec. IV, let us not discard the bottom half of the screen, but rather implement the entire screen with a EMCCD camera that is symmetrically positioned with respect to the dotted horizontal line shown in Fig. 3(a). Let us now equip this camera with two measurement read-out channels, R1 and R2, such that R1 outputs the photocurrent generated due to single-photon-detection events in the top half of the camera screen, and R2 outputs the photocurrent generated due to single-photon-detection events in any part of the camera screen; such a situation can be readily arranged by writing a suitable piece of code with the software that controls the camera. The channel R1 is connected to the amplifier, A and the anharmonic potential controller, thereby providing the probabilistic triggering mechanism required in Fig. 2(b).
In contrast, the channel R2 and the clock signal are inputted to a classical, logical AND circuit, and the output is monitored for a window of time equal to the period of the clock signal. If the output is , then no action is taken; however, if the the output is , then it is assumed that the ancillary photon has been lost, and the entire cycle of quantum operation, which commences with the emission of the ancillary photon and concludes with the formation of , is aborted. A new cycle of quantum operation on a freshly prepared, input, entangled state of modes starts with the next clock cycle.
Importantly, the above classical operation should be carried out only after the formation of the macroscopic superposition of the amplifier and the detector, so as not to interfere with the probabilistic altering of the anharmonicity of the potential. In essence, while the efficiency of conversion from to remains unaltered, the output entangled state of photons becomes significantly error-free, at the cost of reduced speed of operations, which is, fundamentally, determined by the single-photon emission rate.
VII Conclusions and Outlook
Broadly construed, photonic entangled states are described as either mode-mode entangled, or photon-photon entangled, where the relevant quantum subsystems are modes (wave-like) and photons (particle-like), respectively. This classification criterion is also applicable to quantum systems of cold, bosonic atoms, molecules, and ions. Mode-mode entangled states are relatively easier to prepare—for example, by amplitude and beam splitting—than photon-photon entangled states; however, the latter is a key prerequisite for optical quantum information processing. In fact, experimentally constructing photon-photon entangled states is one of the grand challenges of quantum optics.